OpenStudy (anonymous):

Solve by elimination 3x+4y=-1 -9x-4y=13

OpenStudy (anonymous):

3(3x+4y=-1) = (9x+12y=-3) -9x-4y=13 then add x's and y's and number's in the 2 equations and end up with 8y=10 then solve from there.